At a distance of 4,000 ft from the launch site, a spectator is observing a rocket being launched. If the rocket lifts off vertically and is rising at a speed of 600 ft/sec when it is at an altitude of 3,000 ft, how fast is the distance between the rocket and the spectator changing at that instant?

Grade:

2 Answers

Chetan Mandayam Nayakar
12 years ago

distance=s=√(40002+h2), rate of change of distance=ds/dt=(ds/dh)(dh/dt)=(d(√(40002+h2)/dh)(dh/dt),dh/dt=600 ft/sec

archit srivastava
11 years ago

250/3 ft/sec using 2d figure